Find the slope and the y-intercept for the line.
8x - 4y = 12

Answer by usyim88hk(157)

First put this in slope-intercept form:
:
y=mx+b
:
8x - 4y = 12
:
Subtract 8x from both sides
:
8x-4y-8x = 12-8x
-4y = -8x+12
:
Then divide both sides by -4
:
(-4y)/-4 = (-8x+12)/-4
y = -2x+3
:
So now we know that -2 is the slope and 3 is the y-intercept